NOAA Office of Satellite and Product Operations

NOAA 16 DHS Subsystem Summary

Component Status:

Component
Description
Status
AIP1AMSU Information Processor side 1Green
AIP2AMSU Information Processor side 2Green
DTR1ADigital Tape Recorder - 1 A SideGreen
DTR1BDigital Tape Recorder - 1 B SideGreen
DTR2ADigital Tape Recorder - 2 A SideGreen
DTR2BDigital Tape Recorder - 2 B SideGreen
DTR3ADigital Tape Recorder - 3 A SideGreen
DTR3BDigital Tape Recorder - 3 B SideGreen
DTR4ADigital Tape Recorder - 4 A SideGreen
DTR4BDigital Tape Recorder - 4 B SideRed
DTR5ADigital Tape Recorder - 5 A SideRed
DTR5BDigital Tape Recorder - 5 B SideRed
MIRPManipulated Information rate ProcessorGreen
TIP1TIROS Information Processor Side 1Green
TIP2TIROS Information Processor Side 2Green
XSU1Cross-strap Unit Side 1 
XSU2Cross-strap Unit Side 2Green

Subsystem History:

Date Time Component Description
11/09/201000:03:00XSU1Configures STX1 for HRPT.
11/09/201000:03:00XSU1Clears STX1's cross-strap configuration.
11/09/201000:03:00XSU1Clears STX2's cross-strap configuration.
10/09/200716:31:00XSU2Turn on XSU Side 2.
10/09/200716:31:00XSU1Turn off XSU Side 1
11/08/200614:14:40TIP2Returns TIP and all recorders to nominal orbital mode
11/08/200610:50:30DTR4ADwell on analog channel 340 (DTR4 current, NDTR4I)
11/08/200607:26:30DTR3ADwell on analog channel 324 (DTR3 current, NDTR3I)
11/08/200604:03:00DTR2ADwell on analog channel 308 (DTR2 current, NDTR2I)
11/08/200600:38:30DTR1ADwell on analog channel 292 (DTR1 current, NDTR1I)
03/02/200621:15:00XSU1Dwell on analog channel 153 (bus voltage)
03/02/200618:05:00XSU1Dwell on analog channel 457 (SARR 406 side A AGC NSARA4G)
07/19/200517:50:00TIP2SAD Boom dwell test completed. Return to dwelling on bus voltage NPSEBUSV channel number 153.
07/19/200514:45:00TIP2Dwell on NSADBDPY channel number 274 for comparative data in support of N18 pitch ocsillation investigation.
04/18/200515:45:00TIP1Switch from TIP Side 1 to TIP Side 2.
03/29/200413:40:00DTR4ADTR-4A Test
03/24/200414:41:00DTR4BPlayback DTR4B to BOT. End Test
03/24/200414:38:00DTR4BTEST Forward Play DTR4B to EOT
03/24/200412:57:00TIP1TIP back to orbit mode
03/24/200412:55:00TIP1Put TIP in dwell mode, sampling NAVHMTRI
04/21/200300:00:00 TIP Analog telemetry is drifting similar to N15, caused by D-A converter in the TIP.
02/01/200307:30:00DTR4BDTR4B STIP test successful
01/31/200317:27:00DTR4BDTR4B STIP test successful
01/31/200308:20:00DTR4BDTR4B STIP test successful
01/16/200315:30:00DTR4BDTR4B forward and reverse playback test - results negative
01/16/200315:30:00XSU1Cross strap Sband 2 to DTR4
01/14/200316:00:00DTR4BPlayback DTR4B unsuccessful
01/14/200316:00:00XSU1Cross strap SBand 4 to Recorder 4B
01/13/200307:45:00DTR4BDTR4B taken out of operational rotation
01/12/200312:18:00DTR4BG4B locked on B/S only during the PB and no data was recorded on either the "A" or "B" IMUX. Additional playback attempts also showed negative data. 4A nominal. DHS status downgraded to YELLOW.
08/27/200200:00:00TIP1Time Stamp File clock offset reconciliation
06/07/200200:00:00DTR5BDTR5 testing - results negative
06/07/200200:00:00DTR5ADTR5 testing - results negative
05/23/200211:00:00AIP1AIP test to verify ground system compatibility - successful
05/07/200215:00:00TIP1High speed dwell test via macro
12/04/200121:15:00TIP1End high speed dwell testing using macros
12/04/200116:15:00TIP1Begin high speed dwell testing using macros
11/04/200100:00:00DTR3ADTR3A BOT problem has self corrected.
10/10/200115:55:00DTR1BSwapped safestate DTR from 1B to 3A for both processors
10/10/200115:55:00DTR3ASwapped safestate DTR from 1B to 3A for both processors
09/18/200100:00:00TIP1Dwell on channel 449 SAR AGC for 24 hours
09/15/200100:00:00DTR3ADTR3A shows as BOT during playback and record
08/07/200100:00:00TIP1ETCUP to correct spacecraft clock following correction of ground system time measurement software
05/16/200100:00:00TIP1Pointing test with sun sensor dwell for 24 hours
04/05/200100:00:00DTR5ADTR5A record/playback tests - negative results
04/02/200100:00:00DTR5ADTR5A record/playback tests - negative results
03/23/200100:00:00DTR5BDTR5B record/playback tests - negative results
03/18/200100:00:00DTR5BDTR5B record/playback tests - negative results
03/15/200100:00:00DTR5BDTR5B record/playback tests - negative results
02/27/200118:00:00DTR5ADTR5A Boost playback and left at BOT
02/06/200100:00:00DTR5ADTR5 taken out of operational schedule
02/02/200100:00:00DTR5ADTR5A determined to not have recorded boost data.
02/02/200100:00:00DTR5BDTR5B incomplete playback
09/21/200010:56:00DTR2AOn Orbit
09/21/200010:56:00DTR2BOn Orbit
09/21/200010:56:00DTR3BOn Orbit
09/21/200010:56:00DTR4BOn Orbit
09/21/200010:56:00DTR5BOn Orbit
09/21/200010:56:00XSU2On Orbit
09/21/200010:56:00TIP2On Orbit
09/21/200010:56:00AIP1On Orbit
09/21/200010:56:00AIP2On Orbit
09/21/200010:56:00DTR4AOn Orbit
09/21/200010:56:00DTR3AOn Orbit
09/21/200010:56:00DTR1AOn Orbit
09/21/200010:56:00TIP1On Orbit
09/21/200010:56:00MIRPOn Orbit
09/21/200010:56:00XSU1On Orbit
09/21/200010:56:00DTR1BOn Orbit
09/21/200010:56:00DTR5AOn Orbit